find the general solution of the equation x.dy/dx+4y=x⁵.e^x using the method of variation of parameters.


1
Expert's answer
2021-02-28T08:36:02-0500

"x\frac{{dy}}{{dx}} + 4y = {x^5}{e^x}"

consider the corresponding homogeneous equation

"x\frac{{dy}}{{dx}} + 4y = 0 \Rightarrow x\frac{{dy}}{{dx}} = - 4y 
\Rightarrow \frac{{dy}}{y} = - 4\frac{{dx}}{x} \Rightarrow \ln y = - 4\ln x + \ln C = \ln \frac{C}{{{x^4}}} \Rightarrow y = \frac{C}{{{x^4}}}"

Let

"C = C(x)"

Then

"y' = \frac{{C'{x^4} - C \cdot 4{x^3}}}{{{x^8}}} = \frac{{C'x - 4C}}{{{x^5}}}"

substitute in the original equation:

"\begin{array}{l}
x\frac{{C'x - 4C}}{{{x^5}}} + 4 \cdot \frac{C}{{{x^4}}} = {x^5}{e^x}\\
\frac{{C'x}}{{{x^4}}} = {x^5}{e^x}\\
C' = {x^8}{e^x}
\end{array}"

then

"C = \int {{x^8}{e^x}dx} = {x^8}{e^x} - \int {8{x^7}{e^x}dx = } {x^8}{e^x} - 8{x^7}{e^x} + \int {56{x^6}{e^x}} dx="

"={x^8}{e^x} - 8{x^7}{e^x} + \int {56{x^6}{e^x}} dx- \int {336{x^5}{e^x}} dx = {x^8}{e^x} - 8{x^7}{e^x} + 56{x^6} {e^x}- 336{x^5}{e^x} +"

"+ \int {1680{x^4}} {e^x}dx = {x^8}{e^x} - 8{x^7}{e^x} + 56{x^6}{e^x} - 336{x^5}{e^x} + 1680{x^4}{e^x} - \int {6720{x^3}} {e^x}dx ="

"= {x^8}{e^x} - 8{x^7}{e^x} + 56{x^6} {e^x}- 336{x^5}{e^x} + 1680{x^4}{e^x} - 6720{x^3}{e^x} + \int {20160{x^2}{e^x}dx} ="

"={x^8}{e^x} - 8{x^7}{e^x} + 56{x^6} {e^x}- 336{x^5}{e^x} + 1680{x^4}{e^x} - 6720{x^3}{e^x} + 20160{x^2}{e^x} - \int {40320x{e^x}} dx={x^8}{e^x} - 8{x^7}{e^x} + 56{x^6}{e^x} - 336{x^5}{e^x} + 1680{x^4}{e^x} - 6720{x^3}{e^x} + 20160{x^2}{e^x} - 40320x{e^x} + {x^8}{e^x} - 8{x^7}{e^x} + 56{x^6} {e^x}- 336{x^5}{e^x} + 1680{x^4}{e^x} - 6720{x^3}{e^x} + 20160{x^2}{e^x} - 40320x{e^x} + \int {40320{e^x}} dx = {x^8}{e^x} - 8{x^7}{e^x} + 56{x^6}{e^x} - 336{x^5}{e^x} + 1680{x^4}{e^x} - 6720{x^3}{e^x} + 20160{x^2}{e^x} - - 40320x{e^x} + 40320{e^x}+{C_1}"

Then

"\begin{array}{l}
y = \frac{C}{{{x^4}}} = \\
 = \frac{{{e^x}\left( {{x^8} - 8{x^7}e + 56{x^6} - 336{x^5} + 1680{x^4} - 6720{x^3} + 20160{x^2} - 40320x - 40320} \right) + {C_1}}}{{{x^4}}}
\end{array}"